# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

### v5

@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021), and IntentQA (Li et al., 2023a). In addition, to further highlight the strength of our approach on longer videos, we include results on VideoMME’s long splitFu et al. (2024). These datasets are public available and can be used freely for academic research. The first dataset, EgoSchema, consists of 5031 questions and each video lasts three-minute and have multiple choice question. The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA, is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.). The fourth dataset, VideoMME, consists of very long videos—some up to one hour long, with an average duration of 44 minutes, and provides 900 Q&A.
Model Choices & Hyperparameters: For the HKS module, we use the ResNet-18 (He et al., 2016a) for the TSC, CLIP-B/16 (Ranasinghe et al., 2023) for the CKD and GPT-4o for the FKD. We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to GPT-4o, whose model size is expected to be on the scale of 100B-1T. This makes them well-suited for filtering dense frames efficiently. In line with previous state-of-the-art work Wang et al. (2024d); Zhang et al. (2023); Wang et al. (2023), we employ GPT API, especially GPT-4o, for both VLM and LLM. This choice is driven by its cost-effectiveness and lighter computational requirements compared to GPT-4. GPT-4o is used as the VLM for generating captions and as the LLM for answering questions in our framework. We run TSC and CKD on a single NVIDIA RTX A5000, which takes approximately two hours to process 500 questions. We use the default hyperparameters for each vision/language module, as we only perform inference, and set the LLM temperature to 0 to ensure reproducibility. Also, We use single run for our experiments.
@@ 4.2 Evaluation
Model Avg. Frame Captions (Rounded) $\downarrow$
6 7 8 10 11 12 16 32 62 70
VideoAgent 58.4 - 63.2 / 60.2 - 60.8 / 57.4 - - - - -
VideoTree - 61.0 - 61.6 62.2 - - 64.4 66.2 67.0
LVNet (Ours) - - 64.4 - - 68.2 67.8 - - -
Templating Order Acc.
@@
We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 2. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by +7%, is on par with VideoTree while using only 1/5 of the captions, and outperforms TraveLER by +7.8% while utilizing only 12% of the captions.
@@
We next evaluate on the NExT-QA dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet outperforms VideoAgent by +1.6%.
@@
In the IntentQA dataset. LVNet outperforms all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement of +4.8% over the next best model, VideoTree, while using only 13% of the captions (12 vs. 90).
Lastly, Table 4 presents the performance of LVNet on VideoMME’s long split, which consists of videos up to one hour long and compare it to other models using keyframe selection methods. Our method (LVNet) demonstrates strong performance while utilizing only 24 frames, significantly fewer than VideoTree’s 98 frames. LVNet outperforms VideoAgent by +6.0% overall and achieves the highest accuracy in three out of six categories: Knowledge, Artistic Performance, and Multilingual. While VideoTree maintains a slight overall lead, LVNet’s ability to achieve comparable accuracy while processing only one-quarter of the frames highlights its efficiency in handling very long videos. To ensure a fair comparison, all models utilize GPT-4o.
Given the generative nature of VQA tasks as well as the limited availability and noisy nature of fully-annotated video VQA corpora, building generalizable fully-supervised models are challenging for these tasks. Nevertheless, we highlight how our zero-shot and video level training-free framework is competitive with the best supervised approaches on this dataset. This indicates the promise of utilizing pretrained models, especially those equipped with extensive world knowledge and reasoning skills from alternate modality specific learning (i.e. in our cases image domain VLMs and language domain LLMs).
Figure 4: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively. LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer. The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
Qualitative Analysis of Hierarhical Keyframe Selector:
We compare the open-ended responses of LVNet and the uniform sampling method in Figure 4 to understand the effectiveness of the hierarchical keyframe selector in LVNet. The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively. LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames. When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, and measuring tape), which is the correct answer. This leads LVNet to choose the correct option.
@@ 4.3 Ablations
In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS. In all ablations, we use a subset of EgoSchema (Mangalam et al., 2023), composed of 500 videos. Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix A.1
@@
In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp proves more effective than using confidence scores or temporal order alone, as shown in Table LABEL:tab:ablation:visualtemplateorder. In this hybrid approach, high-confidence frames capture short but important segments of the video, while low-confidence keyframes, which are crucial but visually challenging for keyword matching, are temporally ordered to cover broader segments. This hybrid approach outperforms solely temporal ordering and solely confidence-based ordering by +3% and +0.6%, respectively.
Number of Frame Captions:
We performed an ablation study on the number of frame captions, comparing our approach to VideoAgent (Wang et al., 2024b) and VideoTree (Wang et al., 2024e) under similar low caption settings. As shown in Table LABEL:tab:ablation:numberofframecaptions, LVNet achieves the highest accuracy of 68.2% with 12 captions, outperforming VideoAgent (8.4 frames) and VideoTree (12 frames) by +8% and $\sim$+5.7%, respectively. We compare LVNet with VideoAgent+GPT-4o (8.1 frames) and VideoTree+GPT-4o (69.5 frames, $\times$5.8 more), both using GPT-4o for a fair comparison. We take GPT-4o numbers for VideoAgent and VideoTree from Yang et al. (2024) (more details in appendix A.2) and LVNet outperforms them by +5% and +1.2%, respectively.
@@ 5 Conclusion
We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across 3 benchmarks datasets. Our Hierarchical Keyframe Selector demonstrates the effectiveness of keyframe selection in understanding a very long-form video QA. Additionally, we highlight the zero-shot capability for long-form video comprehension of our LVNet framework, which requires no video-level training. Our experiments showcase its significant advantage over previous methods.
@@
  • First, we acknowledge that we are unable to evaluate LVNet and other models with all available VLMs or LLMs due to computational constraints and high costs. However, we carefully select GPT-4o, a state-of-the-art LLM, for our main experiments and provide ablation studies comparing various LLMs (e.g. GPT-3.5, GPT-4, and GPT-4o) to other models to ensure a fair performance comparison, as presented in LABEL:tab:ablation:numberofframecaptions and LABEL:tab:ablation:choiceofllm.
@@
  • Like any LLM-based approach, LVNet is sensitive to prompting. To ensure the transparency, we provide examples of these prompts in Figure 4 and Figure A.7. We also plan to release the code to enable further exploration by other researchers.
@@ Acknowledgements:
This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government [24ZB1200, Research of Human-centered autonomous intelligence system original technology]. This work was also supported by the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 100%)
@@ Appendix A.1 Additional Ablations
Table A.5: Additional ablations experiments on EgoSchema (Mangalam et al., 2023): We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot VQA. Default setting is highlighted.
Choice of LLM:
Table LABEL:tab:ablation:choiceofllm shows that GPT-4o outperforms both GPT-4 and GPT-3.5 by +2.8% and +7.2%, respectively. Given that GPT-4o is more cost-effective and lightweight compared to GPT-4, we have selected it as our default LLM.
@@
Table LABEL:tab:ablation:patchsizeinckd shows the effect of the scales of the patch sizes in the CKD. Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network (Ranasinghe et al., 2023) to match keywords. Our findings show that higher resolutions lead to better accuracy. In LVNet, we use a 14$\times$14 feature map and determine the confidence level of the keyword by selecting the maximum value between the 14$\times$14 patches and the keyword’s text embedding.
Table A.6: Exact (Unrounded) Frame Caption Counts. These values supplement the rounded numbers in Table LABEL:tab:ablation:numberofframecaptions All models are based on either GPT-4o or GPT-4.
Model Avg. Frame Captions $\downarrow$
6.4 7.1 8 8.1 8.4 9.7 10.7 11 11.3 12 16 32 62.4 69.5
VideoAgent 58.4 - - 63.2 60.2 - 60.8 57.4 - - - - - -
VideoTree - 61.0 - - - 61.6 - - 62.2 62.5 66.2 64.4 66.2 67.0
LVNet (Ours) - - 64.4 - - - - - - 68.2 67.8 - - -
Exact Frame Caption Counts:
For completeness, Table A.6 lists the original (unrounded) frame caption counts and corresponding accuracies for VideoAgentWang et al. (2024b), VideoTreeWang et al. (2024e), and LVNet (ours). These values supplement the rounded numbers presented in Table LABEL:tab:ablation:numberofframecaptions of the main text.

### v6

@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema (Mangalam et al., 2023), NExT-QA (Xiao et al., 2021), and IntentQA (Li et al., 2023). In addition, to further highlight the strength of our approach on longer videos, we include results on VideoMME’s long split (Fu et al., 2024). These datasets are public available and can be used freely for academic research. The first dataset, EgoSchema, consists of 5031 questions and each video lasts three-minute and have multiple choice question. The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA, is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.). The fourth dataset, VideoMME, consists of very long videos—some up to one hour long, with an average duration of 44 minutes, and provides 900 Q&A. Collectively, these benchmarks span three orders of magnitude in duration—from short clips ($\sim$44 s on NExT‑QA/IntentQA), to minute‑scale videos (EgoSchema, $\approx$3 min), to hour‑scale footage (VideoMME’s long split, 30–60 min)—allowing us to stress‑test LVNet from seconds to hour‑long videos under a single, training‑free evaluation protocol.
Model Choices & Hyperparameters: For the HKS module, we use the ResNet-18 (He et al., 2016a) for the TSC, CLIP-B/16 (Ranasinghe et al., 2023) for the CKD and GPT-4o for the FKD. We select ResNet-18 and CLIP-B/16 due to their smaller models sizes—0.01B and 0.12B, respectively–which are significantly lighter compared to LLMs that are on a billion parameter scale. In line with previous state-of-the-art work Wang et al. (2024f); Zhang et al. (2023); Wang et al. (2023), we use up to 4 NVIDIA RTX A5000 GPUs and GPT or DeepSeek APIs for running baselines and our setup in all evaluations. Also following prior work, we report results over single evaluation runs.
@@ 4.2 Evaluation
Model Frame Captions
8 12 16
VideoAgent+GPT-4o 63.4 63.6 64.6
VideoTree+GPT-4o 62.8 63.6 66.4
LVNet (GPT-4o) 64.4 68.2 67.8
Templating Order Acc. $\uparrow$
@@
Method LLM AP Frames $\uparrow$ LLM frames $\downarrow$ Acc. $\uparrow$
Qwen-VL 7B 128 128 37.8
Qwen-VL 7B 256 256 OOM
Qwen-VL 7B 1800 1800 OOM
LVNet (DS-V3) 37B 1800 24 53.1
Table 5: Single-Stage Method Comparison: We report Accuracy on VideoMME long split (average video length 41 mins) along with LLM active parameters (LLM AP), total frames processed, and frames input to LLM as tokens / captions (LLM frames). LVNet DeepSeek-V3 (DS-V3) variant is used. Processing lengthy frame sequences with single-stage models at fixed compute becomes infeasible. Inference tested on 4 x 24GB A5000 GPUs. Similar findings in Zeng et al. (2024).
@@
We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 2. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. We reiterate how number of captions input to LLM affects inference compute cost of methods the most.
For EgoSchema(fullset), LVNet achieves 61.1%, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by +7%, is on par with VideoTree while using only 1/5 of the captions, and outperforms TraveLER by +7.8% while utilizing only 12% of the captions.
@@
We next evaluate on NExT-QA dataset, which has a particular focus on both temporal and casual reasoning based question-answer pairs. LVNet achieves state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet outperforms VideoAgent by +1.6%.
@@
In IntentQA dataset, LVNet outperforms all prior work, including de-emphasized models with video-caption pretraining and downplayed models utilizing significantly compute (captions input to LLM) than ours. In fact, LVNet shows a substantial improvement of +4.8% over the next best VideoTree, while using only 13% of captions (12 vs 90).
Lastly, Table 3 shows LVNet’s performance on VideoMME’s long split (Fu et al., 2024), which features videos up to an hour long—significantly exceeding the 12-minute average in MLVU (Zhou et al., 2024) or the 17-minute average in the overall VideoMME. In combination with an open-source LLM (DeepSeek-V3), LVNet outperforms the single-stage keyframe selection methods VideoChat-T (Zeng et al., 2024) and Frame-Voyager (Yu et al., 2024b) by +9.3% and +1.9%, respectively, without any video-level training. Moreover, under a direct head-to-head setting where VideoAgent, VideoTree, and LVNet all use GPT-4o as both the frame captioner and LLM with the same 24-frame budget, LVNet achieves 53.9% accuracy, outperforming VideoAgent(51.3%) and VideoTree(52.3%) by +2.6% and +1.6%, respectively, further confirming its effectiveness on very long videos. See Appendix˜A.3 for the detailed results.
Qualitative Examples:
Due to space, qualitative open-ended responses comparing LVNet to uniform sampling are provided in the Appendix (see Figure˜A.6).
Model FV $\uparrow$ FL $\downarrow$ VC $\downarrow$ TC $\downarrow$ Acc. $\uparrow$
GPT-4o 384 384 \$2.88 $\sim$\$2592 65.3
LVNet (GPT-4o) 1800 24 \$0.19 \$171 53.9
Table 6: API Comparison: We perform cost comparison with using GPT-4o directly on frames vs with LVNet. We report accuracy on VideoMME long split (Acc), frames processed per video (FV), frames / captions input to LLM (FL), per video cost (VC), and total evaluation cost (TC). Our LVNet achieves competitive performance at over 10x less inference cost. GPT-4o accuracy result from official benchmark leaderboard.
4.3 Efficiency & Scalability
[…]
Stages FPS Mem. Frames Lat. (s) Lat. (%)
HKS-1/3 (TSC) 182 19 1800 9.9 16
HKS-2/3 (CKD) 138 9 427 3.1 5
HKS-3/3 (FKD) 1 - 8 9.2 15
Captioning 1 - 24 38.7 62
LLM Q&A - - - 1.0 2
Table 7: LVNet Stage-wise Performance: The table reports per-module elapsed time, FPS, and peak GPU memory for a single VideoMME-long video processed on an RTX-6000 Ada. Within LVNet, the three stage Hierarchical Keyframe Selectors contributes 36 % of the total runtime. It indicates that most compute lies in the captioning and LLM Q&A steps, which are overheads shared by all two-stage pipelines, rather than in the keyframe selection. Mem. and Lat. in the header stands for the GPU memory (GB) and Latency.
[…]
VideoAgent 87.3 2.1
VideoTree 64.2 2.8
LVNet 25.6 7.0
[…]
Model Frames Memory (GB) Inference time (s) FPS
Qwen2.5-VL-72B 8 355 11.8 0.7
Qwen2.5-VL-72B 12 373 16.8 0.7
Qwen2.5-VL-72B 24 OOM OOM OOM
LVNet 1800 19 61.9 29.1
Table 9: Model Efficiency Comparison: End-to-end, LVNet handles 1,800 frames in 61.9 s (29.1 FPS) while using just 19 GB on a single GPU, whereas Qwen2.5-VL-72B requires eight GPUs and 373 GB yet runs $\approx 40\times$ slower at 0.7 FPS. Despite architectural and hardware differences, these figures show that LVNet keyframe selector is lightweight and that the overall system delivers high FPS while remaining single-GPU friendly for long-form video.
Single-stage scaling limits on long videos. Table 5 shows that single-stage models (e.g., Qwen-VL (Bai et al., 2023)) run OOM beyond 128 frames, whereas LVNet achieves higher accuracy with only 24 keyframes passed to the LLM, maintaining tractable compute. It shows the LVNet’s scalability of processing large number of frames with high accuracy.
[…]
Scalability across frame budget and LLM upgrades. Table˜10 summarizes two complementary scaling axes. Left: on VideoMME‑Long, increasing the frame budget from 12 to 24 yields a +3.8% gain. Conversely, on EgoSchema’s 3-min videos, the 12-frame setting is optimal (68.2 %), indicating extra keyframes help LVNet for longer videos. It is consistent with our keyframe‑driven design that benefits from additional informative frames when videos are very long. Right: LVNet improves monotonically with stronger LLMs (GPT‑3.5 $\rightarrow$ GPT‑4 $\rightarrow$ GPT‑4o) without any retraining, highlighting the modularity and scalability of LVNet with newer LLMs.
Frames Acc. (%)
12 50.1
24 53.9   (+3.8)
LLM Acc. (%)
GPT-3.5 61.0
GPT-4 65.4
GPT-4o 68.2
Table 10: Scalability of LVNet. Left: accuracy improves as the frame budget increases for very long videos. Right: newer LLMs yield higher accuracy without retraining, reflecting LVNet’s modularity.
@@ 4.4 Ablations
In this section, we present ablations on key design decisions such as the sorting order in FKD, the number of frames for captions, and the effect of different components in HKS. In all ablations, we use a subset of EgoSchema (Mangalam et al., 2023), composed of 500 videos. Additional ablations about Choice of LLM and Effect of Patch Size on Keyword Matching in CKD are in Appendix˜A.1
Number of Frame Captions:
We conduct an ablation on the number of frame captions, directly comparing our approach with VideoAgent (Wang et al., 2024d) and VideoTree (Wang et al., 2024g) under the same GPT-4o backbone for both frame captioning and LLM reasoning. As shown in Table LABEL:tab:ablation:numberofframecaptions, LVNet achieves the highest accuracy of 68.2% with 12 captions, outperforming both VideoAgent and VideoTree (63.6%) by +4.6% absolute. LVNet also consistently outperforms the baselines at 8 captions (+1.0% / +1.6%) and 16 captions (+3.2% / +1.4%) compared to VideoAgent / VideoTree..
@@
In visual templating, prioritizing frames by keyword confidence scores followed by reordering low-confidence frames based on timestamp proves more effective than using confidence scores or temporal order alone, as shown in Table LABEL:tab:ablation:visualtemplateorder. In this hybrid approach, high-confidence frames capture short but important segments of videos, while low-confidence keyframes, which are crucial but visually challenging for keyword matching, are temporally ordered to cover broader segments. This hybrid approach outperforms solely temporal ordering or confidence-based ordering by +3% and +0.6%, respectively.
@@ 5 Conclusion
We proposed a novel approach for Long-form Video Question Answering (LVQA) that achieves state-of-the-art performance compared to the model using the similar-scale captions across four benchmarks. Our Hierarchical Keyframe Selector demonstrates the effectiveness of keyframe selection in understanding a very long-form video QA. Additionally, we highlight the zero-shot capability for long-form video comprehension of our LVNet framework, which requires no video-level training. Our experiments showcase its significant advantage over existing single-stage and two-stage keyframe selectors.
@@
  • First, we acknowledge that we are unable to evaluate LVNet and other models with all available VLMs or LLMs due to computational constraints and high costs. However, we carefully select GPT-4o and DeepSeek-v3, the LLMs commonly used in video understanding research, for our main experiments and provide ablation studies comparing various LLMs (e.g. GPT-3.5, GPT-4, and GPT-4o) to ensure a fair performance comparison, as presented in LABEL:tab:ablation:numberofframecaptions and LABEL:tab:llm_choice_eff.
@@
  • Like any LLM-based approach, LVNet is sensitive to prompting. To ensure the transparency, we provide examples of these prompts in Figure˜A.6 and Figure˜A.7. We also plan to release the code to enable further exploration by other researchers.
@@ Acknowledgements:
This research was financially supported by the Ministry of Trade, Industry, and Energy (MOTIE), Korea, under the “Global Industrial Technology Cooperation Center(GITCC) program” supervised by the Korea Institute for Advancement of Technology (KIAT).(Task No. P0028420). This research was supported by the National Research Council of Science & Technology(NST) grant by the Korea government(MSIT) (No. GTL25041-000).
@@ Appendix A.1 Additional Ablations
Table A.11: Additional comparisons on EgoSchema subset. Left: uniform 12‑frame sampling (GPT‑4o) vs LVNet at 8 and 12 frames. Right: effect of patch size in CKD.
Uniform Sampling Comparison:
GPT-4o with uniform 12‑frame sampling on the EgoSchema subset scores 62.6%. LVNet outperforms GPT-4o, reaching 64.4% with just 8 frames and 68.2% with 12 frames (Table LABEL:tab:appendix_uniform). To keep the evaluation both methodical and cost‑effective, we first ran GPT-4o and LVNet on the EgoSchema subset as validation to find the optimal settings, then transferred those settings to the full EgoSchema test set and the remaining benchmarks. This procedure avoids the $\mathdollar 3{,}000$ API expense of running GPT-4o uniformly across all four datasets while still demonstrating LVNet’s stronger generalization. For LLM choices, see Sec. 4.3.
@@
Table LABEL:tab:ablation:patchsizeinckd shows the effect of the scales of the patch sizes in the CKD. Since keywords can represent activities spanning the entire image or confined to a small region, we adjust the resolution of the visual feature map output from the spatially aware contrastive image pre-training (CLIP) network (Ranasinghe et al., 2023) to match keywords. Our findings show that higher resolutions lead to better accuracy. In LVNet, we use a $\times14$ feature map and determine the confidence level of the keyword by selecting the maximum value between the $\times14$ patches and the keyword’s text embedding.